OpenStudy (anonymous):

Write the recursive formula for the geometric sequence. a1 = -2 a2 = 8 a3 = -32

OpenStudy (anonymous):

look the difference between -2 and 8 \[−2^{1}=−2\] \[−2^{1+n}=8\]

OpenStudy (anonymous):

i guess you can figure out

OpenStudy (anonymous):

noo

OpenStudy (anonymous):

an = -4 • an-1 ?

OpenStudy (anonymous):

\[-2^{1+(2 \times 0)}=-2, -2^{1+(2 \times 1)}=8, -2^{1+(2 \times 2)}=-32\]